### v1

# RED-DiffEq: Regularization by denoising diffusion models for solving inverse PDE problems with application to full waveform inversion

Abstract
Partial differential equation (PDE)-governed inverse problems are fundamental across various scientific and engineering applications; yet they face significant challenges due to nonlinearity, ill-posedness, and sensitivity to noise. Here, we introduce a new computational framework, RED-DiffEq, by integrating physics-driven inversion and data-driven learning. RED-DiffEq leverages pretrained diffusion models as a regularization mechanism for PDE-governed inverse problems. We apply RED-DiffEq to solve the full waveform inversion problem in geophysics, a challenging seismic imaging technique that seeks to reconstruct high-resolution subsurface velocity models from seismic measurement data. Our method shows enhanced accuracy and robustness compared to conventional methods. Additionally, it exhibits strong generalization ability to more complex velocity models that the diffusion model is not trained on. Our framework can also be directly applied to diverse PDE-governed inverse problems.
Keywords:
partial differential equation; inverse problem; full waveform inversion; regularization; diffusion model
1 Introduction
Inverse problems governed by partial differential equations (PDEs) play a crucial role in diverse scientific and engineering domains [1, 2, 3, 4], from medical imaging [5] and fluid dynamics [6] to geophysical studies [7]. These problems involve inferring unknown parameters or fields from indirect observations, where the forward process is described by PDEs. A prominent example is full waveform inversion (FWI) in seismic imaging, which aims to infer subsurface structures for resource exploration, environmental studies, and seismic hazard assessment [8]. Unlike linear inverse problems, PDE-governed inverse problems usually formulate the inference as a nonconvex and nonlinear optimization task, iteratively minimizing the mismatch between observed data and current PDE solution to optimize unknown parameters [9].
These inverse problems are intrinsically challenging due to their nonlinearity and ill-posedness, making them highly sensitive to stochastic perturbations, such as measurement noise and missing data [9, 10, 11, 12]. In FWI, for instance, these challenges manifest as cycle skipping and convergence to local minima, particularly in complex settings or with limited data quality [13]. Various regularization strategies have been proposed to address these issues. Classical approaches, such as Tikhonov regularization, promote solution smoothness [14], while total variation (TV) regularization preserves sharp boundaries [15, 16]. These regularization techniques enhance the optimization process, improve convergence, and mitigate the effects of ill-posedness. They remain state-of-the-art regularization methods in full waveform inversion [17].
Machine learning has introduced innovative approaches to PDE-governed inverse problems. Deep learning models, particularly convolutional neural networks, have demonstrated success in learning direct mappings from observations to underlying parameters in synthetic environments [18, 19, 20]. These approaches offer potential advantages in computational efficiency and eliminate sensitivity to initial parameters [21]. However, their practical application is often limited by poor generalization, especially when handling noise, missing data, or scenarios not represented in training datasets [22, 23, 24]. To address these limitations, advanced physics-informed machine learning frameworks have emerged. Physics-informed neural networks (PINNs) integrate physical constraints directly into the learning process, demonstrating promising results for solving inverse problems [25, 26, 27, 28, 29, 30, 31, 32]. This integration enhances solution consistency with known physical laws while maintaining computational efficiency [33, 34, 35, 36]. Building on this foundation, deep neural operators [37, 38, 39, 40, 41] have advanced the field by learning sophisticated mappings between function spaces, offering improved generalization across diverse physical scenarios [42, 43, 44, 45, 46, 47]. Despite these significant advances, both approaches continue to face challenges with out-of-distribution data, particularly when confronted with complex real-world problems.
Another promising recent development in machine learning has been the emergence of diffusion models, which have revolutionized generative modeling through their iterative denoising approach to data generation, effectively capturing complex probability distributions [48, 49, 50]. These models offer unique advantages for inverse problems due to their ability to incorporate prior information and their inherent robustness to noise [51, 52, 53, 54]. Diffusion models have also demonstrated success in reconstructing complex fields from limited or noisy observations in various domains [55, 56, 57]. However, their application to inverse problems involving explicit PDE solvers as forward operators remains under-explored. In FWI, for instance, initial efforts have primarily focused on leveraging diffusion models as prior regularizers [58, 59, 60], or leveraging a joint diffusion architecture to solve the velocity model in the latent space [61]. These approaches, while promising in synthetic scenarios, have not yet always outperformed conventional regularization techniques like Tikhonov regularization.
Here, we develop a general framework, regularization by denoising using diffusion models for partial differential equations (RED-DiffEq), that integrates diffusion models directly into PDE-governed inverse problems as a regularization mechanism. Our approach leverages diffusion models to learn robust prior distributions over plausible solutions from synthetic datasets. By employing the pretrained diffusion model as a physics-aware regularization term, the regularization effect is achieved by calculating the residual between actual and predicted noise.
We demonstrate the effectiveness of RED-DiffEq through extensive validation on FWI, a challenging inverse problem in geophysics. Our results consistently show superior performance in accuracy and robustness compared to conventional regularization methods and existing diffusion-based approaches, both quantitatively and qualitatively. Notably, when trained on the OpenFWI dataset [62], RED-DiffEq demonstrates strong generalization ability to benchmark models such as Marmousi [63] and Overthrust [64]. This highlights its potential for application to real-world field data.
2 Results
We first introduce full waveform inversion in Section 2.1 and present the RED-DiffEq framework in Section 2.2. We then evaluate the effectiveness of RED-DiffEq in three scenarios from the OpenFWI benchmark (Section 2.3). We also test RED-DiffEq on the Marmousi (Section 2.4) and Overthrust (Section 2.5) data to demonstrate its generalizability.
2.1 Full waveform inversion
Full waveform inversion seeks to reconstruct high-resolution subsurface velocity models from seismic measurement data (potentially missing or noisy) by leveraging the physics of wave propagation (Fig. 1). In the acoustic approximation, wave propagation is governed by
$\frac{1}{\mathbf{x}^{2}(\mathbf{r})}\frac{\partial^{2}\mathbf{u}(\mathbf{r},t)}{\partial t^{2}}\;-\;\nabla^{2}\mathbf{u}(\mathbf{r},t)\;=\;q(\mathbf{r},t),$
where $\mathbf{r}$ represents the spatial coordinate, $t$ is time, $\nabla^{2}$ is the spatial Laplacian, and $\mathbf{x}(\mathbf{r})$ denotes the subsurface velocity model, $\mathbf{u}(\mathbf{r},t)$ is the seismic wavefield, and $q(\mathbf{r},t)$ represents the source term. The process of solving this equation for a given velocity model to predict the wavefield is known as forward modeling.
Traditional physics-driven numerical methods for FWI are formulated as an optimization problem that iteratively adjusts the velocity model x to minimize an objective function, which reconciles simulated data with observed data:
$\underset{\mathbf{x}}{\text{argmin}}\ \|\mathbf{u}_{\text{data}}-f_{\text{PDE}}(\mathbf{x})\|^{2}_{2}+\lambda\cdot R(\mathbf{x}).$ (1)
Here, the first term measures the misfit between the observed seismic data, $\mathbf{u}_{\text{data}}$, and the simulated data, $f_{\text{PDE}}(\mathbf{x})$. The second term is a regularization term weighted by $\lambda$, which is crucial for constraining the solution to be geologically plausible and for improving the convergence of the inversion.
While traditional regularization methods, such as Tikhonov and TV, improve performance, they often fail to produce geologically realistic models. Tikhonov regularization, for instance, tends to introduce over-smoothing, while TV can create “staircase” artifacts. Moreover, this physics-driven approach leads to an ill-posed inverse problem that is highly sensitive to the choice of the initial velocity model and the presence of measurement noise. For a more detailed description of FWI, please refer to Section S1.
Figure 1: Forward modeling and full waveform inversion. The velocity model (top) with five sources (A–E) generates seismic data under different conditions: clean data (noise-free), Gaussian noise, and missing traces.
2.2 RED-DiffEq for full waveform inversion
To solve inverse problems of PDEs like FWI, we develop a new regularization framework named RED-DiffEq (Algorithm S1) based on the regularization by denoising (RED) method [65]. Let $p_{\text{data}}$ denote the density of the prior distribution over clean, geologically plausible velocity models. For a velocity model $\mathbf{x}$ at iteration $k$ (denoted $\mathbf{x}_{k}$), the RED regularizer is
$R(\mathbf{x}_{k})\;=\;\mathbb{E}_{t,\boldsymbol{\epsilon}}[\mathbf{x}_{k}^{\top}\!\bigl(\mathbf{x}_{k}-\mathcal{D}_{\theta}(\mathbf{x}_{k};t,\boldsymbol{\epsilon})\bigr)].$ (2)
$\mathcal{D}_{\theta}$ is a Tweedie-inspired denoise operator that nudges $\mathbf{x}_{k}$ toward regions of higher density. To define $\mathcal{D}_{\theta}$, let $\gamma(t)\!\in\!(0,1]$ be a noise schedule (Algorithm S2), where $t\sim\mathcal{U}\{1{:}T\}$, and $T$ is a hyperparameter. In each iteration of RED-DiffEq, we first compute a noisy velocity model $\mathbf{x}_{k,t}$ from the Variance-Preserving (VP) corruption function by
$\mathbf{x}_{k,t}\;=\;\sqrt{\gamma(t)}\,\mathbf{x}_{k}\;+\;\sqrt{1-\gamma(t)}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I}).$ (3)
Then, by Tweedie’s identity for Gaussian channels, we set
$\mathcal{D}_{\theta}(\mathbf{x}_{k};t,\boldsymbol{\epsilon})\;=\;\frac{1}{\sqrt{\gamma(t)}}\Bigl(\mathbf{x}_{k,t}\;+\;\bigl(1-\gamma(t)\bigr)\,s_{\text{data},t}(\mathbf{x}_{k,t})\Bigr).$ (4)
In particular, $s_{\text{data},t}(\mathbf{z})=\nabla_{\mathbf{z}}\log p_{\text{data},t}(\mathbf{z})$, where $p_{\text{data},t}$ is the noisy marginal density obtained by applying the VP corruption in Eq. (3) to samples from the distribution with the density $p_{\text{data}}$.
We approximate this score function by a neural network $\hat{\boldsymbol{\epsilon}}_{\theta}$ (typically a U-Net as in Fig. 2a) using the Denoising Diffusion Probabilistic Model (DDPM) on a corpus of clean velocity models $\mathbf{x}_{\text{data}}$ (Fig. 2b). After training, we estimate the score in Eq. (4) as
$s_{\text{data},t}(\mathbf{x}_{k,t})\;\approx\;-\frac{\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)}{\sqrt{1-\gamma(t)}}.$ (5)
Then substituting Eq. (5) into Eq. (4) yields the practical denoise operator
$\mathcal{D}_{\theta}(\mathbf{x}_{k};t,\boldsymbol{\epsilon})\;=\;\frac{1}{\sqrt{\gamma(t)}}\Bigl(\mathbf{x}_{k,t}\;-\;\sqrt{1-\gamma(t)}\,\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)\Bigr).$ (6)
Finally, substituting Eq. (6) into Eq. (2) gives the RED regularizer by
$R(\mathbf{x}_{k})\;=\;\mathbb{E}_{t,\boldsymbol{\epsilon}}\left[w(t)\,\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr)\right],\quad w(t)\;=\;\sqrt{\tfrac{1-\gamma(t)}{\gamma(t)}},$
where $\boldsymbol{\epsilon}$ is the random Gaussian noise we draw in Eq. (3). In our experiments we adopt a pragmatic variant that drops $w(t)$ and uses a constant $\lambda$, which empirically improves stability and convergence. Therefore, the empirical inversion objective minimized at iteration $k$ is given by
$\mathcal{L}(\mathbf{x}_{k})\;=\;\bigl\|\,\mathbf{u}_{\text{data}}-f_{\mathrm{PDE}}(\mathbf{x}_{k})\,\bigr\|_{2}^{2}\;+\;\lambda\,\underbrace{\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr)}_{\hat{R}(\mathbf{x}_{k})},$
where $\hat{R}$ provides an unbiased Monte Carlo (MC) estimator of the RED regularizer (Fig. 3). In practice, we randomly sample $(t,\boldsymbol{\epsilon})$ at each iteration, which injects stochasticity into the inversion process. We present the details of RED-DiffEq in Sec. 4.1.
Figure 2: Diffusion model architecture and generated samples of velocity map. (a) Schematic illustration of the U-Net architecture with time embedding used in the diffusion model. (b) Overview of the complete diffusion process, showing both the forward noising process and the learned reverse denoising process. (c) Comparison between velocity maps from the training dataset (top row) and unconditionally generated velocity maps from a single pretrained diffusion model (bottom row).
Figure 3: Schematic illustration of RED-DiffEq for full waveform inversion. (a) An overview of the inversion process that iteratively updates the velocity model. (b) An illustration of each iteration step of RED-DiffEq. (c) Calculation of the diffusion-based regularization term. (d) Post-processing refinement step using the pretrained diffusion model to further refine the velocity map after the main inversion process is finished.
2.3 Validation on the OpenFWI benchmark
We conducted a comprehensive evaluation of RED-DiffEq on the OpenFWI benchmark [62] under three distinct scenarios: clean seismic data, seismic data with Gaussian noise contamination, and seismic data with missing traces. The OpenFWI dataset has a collection of numerical solutions to the acoustic wave equation representing diverse geological structures, providing a comprehensive benchmark for evaluating seismic inversion methods. For our experiments, we selected four particularly challenging families from this dataset: Curve Fault (CF-B), Flat Velocity (FV-B), Flat Fault (FF-B), and Curve Velocity (CV-B) (see some examples in Fig. 2c). Each family represents distinct geological structures that are representative of complex subsurface formations encountered in real-world scenarios. Specifically, CF-B features intricate curved fault systems with varying angles and intersections. FV-B consists of horizontally layered velocity structures. FF-B contains straight fault lines cutting through the model. Lastly, CV-B exhibits curved velocity variations with smooth transitions between layers. For benchmarking, we compared our method against established physics-driven approaches, including standard FWI (no regularization), FWI with Tikhonov regularization, FWI with total variation (TV) regularization, and a recent diffusion model based method [58].
2.3.1 Clean seismic data
We first evaluated various methods on clean seismic data using 100 previously unseen samples from each of the four chosen geological families in the OpenFWI dataset, and all the quantitative metrics are the average of these 400 test cases. Each sample was optimized for 300 iterations, starting from an initial model obtained by applying a Gaussian filter (with a standard deviation $\sigma=10$) to the ground-truth velocity model. Fig. 4a shows representative velocity inversion examples across the four different geological families. The unregularized standard FWI method produces a large error and fails to capture essential geological features. While Tikhonov regularization improves upon the standard method, it produces an overly smoothed model that loses critical details at layer boundaries. TV preserves sharp edges but introduces “staircase” artifacts that distort the geology. The diffusion-based approach by Wang et al. shows promise with simple structures but struggles with fine-scale fault structures. Our RED‑DiffEq framework outperforms all baselines across the four geological families, simultaneously preserving high‑contrast velocity discontinuities, such as fault planes, while maintaining the smooth lateral and vertical velocity gradients that characterize individual strata. RED‑DiffEq produces the most faithful reconstructions, even for the complex CV‑B models (Fig. 4a, last row).
Figure 4: Results on the OpenFWI dataset. (a)Qualitative comparison among different regularization methods. (b) Quantitative comparison among different regularization methods through 300 iterations. (c) Vertical profile results of different regularization methods. (d) Performance of different regularization methods on different noise levels. (e) Performance of different regularization methods on different number of missing traces.
The convergence analysis (Fig. 4b, top row) reveals that RED-DiffEq consistently achieves the lowest RMSE and MAE, and highest SSIM throughout the optimization process. Statistical analysis using raincloud plots (Fig. 4b, bottom row) further confirms RED-DiffEq’s superior performance, revealed by narrower error distributions centered on better median values than other methods. Moreover, the more compact interquartile range indicates greater consistency and stability across diverse geological scenarios.
We further evaluated the vertical profiles of the selected inversion examples (Fig. 4a, top row) by extracting velocity traces at the midpoint (350 m) of the velocity map. The resulting profiles (Fig. 4c) demonstrate that RED-DiffEq outperforms the other methods by more closely matching the ground truth. Notably, even in depth intervals where all methods perform reasonably well (e.g., from 0 m to 300 m), other methods exhibit oscillations around the ground truth, whereas RED-DiffEq effectively suppresses these artifacts.
2.3.2 Noisy seismic data
We evaluated the robustness of various inversion methods by adding Gaussian noise with standard deviations from 0.1 to 0.5 to the seismic data. The comparative results reveal distinct behaviors across different approaches under noisy conditions (Fig. 4d).
RED-DiffEq exhibits superior noise resilience, maintaining sharp geological boundaries and coherent structural features at all noise levels. Most notably, it accurately reconstructs the distinct layered structures and preserves the continuous interfaces that are evident in the ground truth (Fig. 4d, right column). In contrast, the unregularized FWI approach shows significant noise-induced artifacts, particularly in the deeper regions of the velocity model, with increasing deterioration at higher noise levels. Tikhonov regularization reduces noise but introduces excessive smoothing, blurring the sharp boundaries between layers present in the ground truth. While total variation regularization preserves some sharp boundaries, it produces notable staircase artifacts that erase finer details, especially when the standard deviation of noise is 0.5, which further hinders the geological interpretability of the result. The method by Wang et al. shows moderate improvement over traditional approaches but struggles to maintain structural coherence under high noise conditions, exhibiting patchy artifacts in the reconstructed velocity model. By comparison, RED-DiffEq achieves the best balance between noise suppression and feature preservation, maintaining geological plausibility and interpretability even under severe noise contamination.
2.3.3 Seismic data with missing traces
We evaluated the performance of all methods in missing trace scenarios by randomly removing 5 to 20 traces from the seismic data. All methods exhibit solid robustness (Fig. 4e), with only gradual performance degradation as the number of missing traces increases. This behavior highlights the physics-driven resilience of FWI, as further supported by the two-way ANOVA test in Section S5, where the main effect of missing traces is not statistically significant ($p>0.05$).
However, visual examination of the reconstructed velocity models reveals notable differences in reconstruction quality (Fig. 4e, right column). RED-DiffEq maintains the highest fidelity of geological structures, successfully preserving all the essential features in the upper and bottom regions of the velocity model. In contrast, all other methods, while stable in metrics, shows visible deterioration in resolving deeper structures as missing traces increase.
These results highlight an important distinction between quantitative metrics and qualitative geological interpretability in FWI. While missing traces may not significantly impact numerical accuracy, the ability to maintain geologically meaningful reconstructions becomes a crucial differentiator among methods.
2.4 Validation on the Marmousi benchmark
To further evaluate our method’s performance, we conducted experiments on the widely used Marmousi model [63], which features complex geological structures, sharp velocity contrasts, and intricate subsurface heterogeneity. This challenging benchmark differs substantially from the OpenFWI dataset used to train our diffusion model, thus providing a rigorous test of out-of-distribution performance.
We initialized the velocity model by applying a Gaussian filter (with $\sigma$ ranging from 20 to 30) to the ground-truth model, effectively removing fine-scale details to mimic practical scenarios characterized by limited prior information. RED-DiffEq accurately reconstructs critical geological features within 300 iterations, recovering complex fault systems, sharp velocity discontinuities, and layered strata (Fig. 5a). Crucially, it maintains this high-fidelity performance across all considered initial conditions, as evidenced by the consistently superior results in RMSE, MAE, and SSIM (Fig. 5b). While these metrics do show a mild sensitivity to initialization at higher $\sigma$, the degradation remains minimal compared to that of benchmark methods. In contrast, unregularized physics-driven methods exhibit high-frequency artifacts, Tikhonov regularization yields only marginal improvements, TV regularization suffers from staircase effects, and Wang et al.’s diffusion-based technique shows limited generalization capacity, performing even worse than the physics-driven baseline in these more complex settings.
Figure 5: Results on the Marmousi model. (a) Qualitative comparisons among different regularization methods. (b) Quantitative performance of each method under various initial parameters. The variable $\sigma$ indicates the standard deviation of the Gaussian filter applied to the ground-truth data to generate the initial model. Larger $\sigma$ implies less preserved information. (c) Vertical velocity profiles at 4.9 km position comparing the inversion results to the ground truth.
To examine velocity reconstructions in greater detail, we compare vertical profiles at a position of 4.9 km for Marmousi (Fig. 5c). RED-DiffEq closely follows the ground truth across diverse depth intervals, capturing both sharp velocity transitions and more gradual gradients. Other methods deviate substantially at mid-range to deeper sections, underscoring their difficulties in preserving critical geological features under limited initial information.
2.5 Validation on the Overthrust benchmark
We extended our evaluation to the Overthrust model [64], which also presents complex geological structures, sharp velocity contrasts, and intricate subsurface heterogeneity. This model serves as another challenging benchmark that differs from our training data, further testing our method’s generalization capabilities.
Similar to the Marmousi experiments, we initialized the velocity model using a Gaussian filter with $\sigma$ ranging from 20 to 30 applied to the ground-truth model. RED-DiffEq successfully reconstructs the essential geological features within 300 iterations (Fig. 6a). The quantitative results in Fig. 6b confirm that our method consistently outperforms benchmark methods across all initial conditions in terms of RMSE, MAE, and SSIM. While there is some performance degradation at higher $\sigma$ values, RED-DiffEq remains significantly more robust than other methods.
Figure 6: Results on the Overthrust model. (a) Qualitative comparisons among different regularization methods. (b) Quantitative performance (RMSE, MAE, and SSIM) of each method under various initial parameters. The variable $\sigma$ indicates the standard deviation of the Gaussian filter applied to the ground-truth data to generate the initial model. Larger $\sigma$ implies less preserved information. (c) Vertical velocity profiles at 3.5 km position comparing the inversion results to the ground truth.
The vertical velocity profile at 3.5 km for the Overthrust model (Fig. 6c) demonstrates RED-DiffEq’s ability to accurately capture velocity variations across different depths. Other methods struggle with accurately recovering the velocity structure, particularly in deeper sections. RED-DiffEq’s performance on the Overthrust model further confirms its ability to generalize to different geological settings while maintaining high reconstruction fidelity, demonstrating its practical value for real-world seismic imaging applications.
3 Conclusions
We introduce RED-DiffEq, a framework that integrates diffusion models as a regularization method for solving inverse PDE problems. Through comprehensive experiments on FWI, we have demonstrated the key advantages of RED-DiffEq. RED-DiffEq not only performs better than other methods on clean seismic data, but also shows remarkable robustness against two types of data imperfections: Gaussian noise contamination and missing traces, where conventional methods typically fail or require careful parameter tuning. RED-DiffEq also demonstrates strong generalization capabilities beyond its training data distribution. Despite being trained only on the OpenFWI dataset, it successfully reconstructs complex geological structures in the challenging Marmousi and Overthrust benchmarks.
The RED-DiffEq framework is inherently general and could be easily adapted to other PDE-governed inverse problems. RED-DiffEq represents a step forward in combining data-driven deep learning with physics-based inversion, offering a promising direction for solving challenging inverse problems while maintaining robust performance and physical consistency. Future research directions include incorporating physics-informed constraints in the diffusion model training and exploring applications to other inverse problems.
4 Methods
4.1 RED-DiffEq framework
We present RED-DiffEq as a two-stage framework that integrates a learned prior into a PDE-governed inverse problem via regularization by denoising (RED). The method comprises two stages: (1) a pretraining stage that learns a generative prior and provides an approximation to the marginal score of the noisy data distribution (Fig. 2), and (2) an inversion stage that uses a Tweedie-inspired denoise operator as a regularizer (Fig. 3). In the RED regularizer (Eq. (2)), the operator $\mathcal{D}_{\theta}$ encourages the current solution $\mathbf{x}_{k}$ toward regions of higher plausibility under the prior distribution with density $p_{\text{data}}$.
Stage 1: DDPM pretraining.
We train a denoising diffusion probabilistic model to predict the Variance-Preserving (VP) corruption noise, which yields an estimator of the marginal score of the noisy prior under the standard $\epsilon$-prediction parameterization (Fig. 2a). The VP forward corruption is
$\mathbf{x}_{\text{data},t}\;=\;\sqrt{\gamma(t)}\,\mathbf{x}_{\text{data}}\;+\;\sqrt{1-\gamma(t)}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\quad\gamma(t)\in(0,1],$ (7)
which induces the conditional density
$q\!\big(\mathbf{x}_{\text{data},t}\mid\mathbf{x}_{\text{data}}\big)\;=\;\mathcal{N}\!\Big(\mathbf{x}_{\text{data},t};\ \sqrt{\gamma(t)}\,\mathbf{x}_{\text{data}},\ (1-\gamma(t))\,\mathbf{I}\Big),$
and the noisy marginal
$p_{\text{data},t}\!\big(\mathbf{x}_{\text{data},t}\big)=\!\int p_{\text{data}}\!\big(\mathbf{x}_{\text{data}}\big)\,q\!\big(\mathbf{x}_{\text{data},t}\mid\mathbf{x}_{\text{data}}\big)\,d\mathbf{x}_{\text{data}}.$ (8)
We train a U-Net $\hat{\boldsymbol{\epsilon}}_{\theta}(\cdot,t)$ with the standard $\epsilon$-prediction loss:
$\mathcal{L}_{\text{DDPM}}(\theta)\;=\;\mathbb{E}_{\,t\sim\mathcal{U}\{1:T\},\,\mathbf{x}_{\text{data}}\sim P_{\text{data}},\,\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\!\left[\left\|\,\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{\theta}\!\big(\mathbf{x}_{\text{data},t},\,t\big)\,\right\|_{2}^{2}\right].$
Under this parameterization, the score of the noisy marginal (Eq. (8)) given an arbitrary noisy sample $\mathbf{z}$ is approximated by
$s_{\text{data},t}\!\big(\mathbf{z}\big)\;=\;\nabla_{\mathbf{z}}\log p_{\text{data},t}\!\big(\mathbf{z}\big)\;\approx\;-\,\frac{\hat{\boldsymbol{\epsilon}}_{\theta}\!\big(\mathbf{z},t\big)}{\sqrt{1-\gamma(t)}},$ (9)
and we provide more details in Sec. S3.1.
Stage 2: Tweedie-inspired denoise operator and inversion.
During inversion, the prior network is frozen and we use Tweedie’s identity for the Gaussian (VP) channel to define the denoise operator. For a current solution $\mathbf{x}_{k}$, we sample $t\!\sim\!\mathcal{U}\{1:T\}$ and $\boldsymbol{\epsilon}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$, compute $\mathbf{x}_{k,t}$ with $\gamma(t)\in(0,1]$ using Eq. (3), and apply Tweedie’s identity (Eq. (4)) to obtain (see Sec. S3.2 for more details)
$\mathcal{D}_{\theta}(\mathbf{x}_{k},t)\;=\;\frac{1}{\sqrt{\gamma(t)}}\Bigl(\mathbf{x}_{k,t}+\bigl(1-\gamma(t)\bigr)\,s_{\text{data},t}(\mathbf{x}_{k,t})\Bigr).$ (10)
Using the $\epsilon$-prediction surrogate from Eq. (9), we evaluate Eq. (10) in practice as
$\mathcal{D}_{\theta}(\mathbf{x}_{k},t)\;=\;\frac{1}{\sqrt{\gamma(t)}}\Bigl(\mathbf{x}_{k,t}-\sqrt{1-\gamma(t)}\,\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)\Bigr).$ (11)
Plugging Eq. (11) into the RED regularizer (Eq. (2)) and expanding with Eq. (3) yields (see Sec. S3.3)
$R(\mathbf{x}_{k})\;=\;\mathbb{E}_{t,\boldsymbol{\epsilon}}\left[w(t)\,\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr)\right],\quad w(t)\;=\;\sqrt{\frac{1-\gamma(t)}{\gamma(t)}}.$ (12)
During experiments, we use a single-sample Monte Carlo estimate of the expectation. Moreover, we empirically find that dropping $w(t)$ yields better convergence and performance. Therefore, we fix the regularization weight empirically as a constant $\lambda$.
Our final inversion objective at iteration $k$ combines the data fidelity (PDE misfit) with the RED regularization term:
$\mathcal{L}(\mathbf{x}_{k})\;=\;\bigl\|\,\mathbf{u}_{\text{data}}-f_{\mathrm{PDE}}(\mathbf{x}_{k})\,\bigr\|_{2}^{2}\;+\;\lambda\;\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr).$ (13)
We minimize Eq. (13) with respect to $\mathbf{x}_{k}$ using gradient descent optimization. To reduce the computational cost, we stop gradients through the denoiser for computing $\nabla_{\mathbf{x}}R(\mathbf{x}_{k})$, i.e., we treat $\hat{\boldsymbol{\epsilon}}_{\theta}$ as a constant with respect to $\mathbf{x}_{k}$ in the regularizer term, whose performance has been empirically validated in Refs. [57, 66]. At each iteration, we resample $(t,\boldsymbol{\epsilon})$, which introduces stochasticity to the optimization.
4.2 Evaluation metrics
We assess reconstruction quality using three common metrics: the root mean square error (RMSE), the mean absolute error (MAE), and the structural similarity index measure (SSIM). Let $\{\mathbf{x}_{i}\}_{i=1}^{N}$ and $\{\hat{\mathbf{x}}_{i}\}_{i=1}^{N}$ denote the ground-truth and recovered models across a spatial domain of $N$ points, respectively. The RMSE and MAE are defined by
RMSE $=\sqrt{\frac{1}{N}\,\sum_{i=1}^{N}\bigl(\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\bigr)^{2}},$
MAE $=\frac{1}{N}\,\sum_{i=1}^{N}\bigl|\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\bigr|.$
Both measure point-wise differences, with RMSE emphasizing larger errors and MAE capturing average deviations.
SSIM [67] compares structural and perceptual similarities between $\mathbf{x}_{i}$ and $\hat{\mathbf{x}}_{i}$. Its value ranges from 0 to 1, where higher scores indicate greater visual and structural fidelity. Formally,
$\text{SSIM}(\mathbf{x},\hat{\mathbf{x}})\;=\;\frac{\bigl(2\mu_{x}\mu_{\hat{x}}+C_{1}\bigr)\bigl(2\sigma_{x\hat{x}}+C_{2}\bigr)}{\bigl(\mu_{x}^{2}+\mu_{\hat{x}}^{2}+C_{1}\bigr)\bigl(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}+C_{2}\bigr)},$
where $\mu_{x},\mu_{\hat{x}}$ are the local means of $\mathbf{x}$, $\hat{\mathbf{x}}$, $\sigma_{x}^{2}$, $\sigma_{\hat{x}}^{2}$ are their variances, $\sigma_{x\hat{x}}$ is the covariance, and $C_{1},C_{2}$ are small constants for numerical stability. Note that the local statistics are computed over small sliding windows (e.g., $11\times 11$) centered at each pixel location. The resulting local SSIM values are then averaged over the entire image to yield the final score. While RMSE and MAE quantify overall intensity errors, SSIM captures perceptual and structural consistency, providing complementary insights into the quality of the reconstructed velocity models.
Acknowledgments
This work was supported by the U.S. Department of Energy Office of Advanced Scientific Computing Research under Grants No. DE-SC0025593 and DE-SC0025592.
Author contributions
All authors contributed to the conceptualization and design of the study. S.S. and M.Z. implemented the code and conducted the experiments. All authors took part in writing and revising the manuscript. Y.L. and L.L. supervised the research effort, providing guidance and critical feedback throughout.
Competing interests
The authors declare no competing interests.
Code availability
The codes in this study will be publicly available at the GitHub repository https://github.com/lu-group/red-diffeq.
Appendix S1 Acoustic full waveform inversion
Full waveform inversion (FWI) is an advanced seismic imaging technique that reconstructs high-resolution subsurface velocity models by utilizing the full seismic waveform. Unlike conventional methods that depend on simplified assumptions or isolated wave arrivals, FWI leverages the full physics of seismic wave propagation by directly embedding the governing equations—typically acoustic or elastic wave equations—into the inversion framework.
In the acoustic approximation, commonly used in exploration geophysics, wave propagation is governed by the following second-order PDE:
$\frac{1}{\mathbf{x}^{2}(\mathbf{r})}\frac{\partial^{2}\mathbf{u}(\mathbf{r},t)}{\partial t^{2}}-\nabla^{2}\mathbf{u}(\mathbf{r},t)=q(\mathbf{r},t),$
where $\mathbf{u}(\mathbf{r},t)$ represents the seismic wavefield, $\mathbf{x}(\mathbf{r})$ denotes the spatial distribution of the acoustic velocity model, $q(\mathbf{r},t)$ is the seismic source term (located at points A through E in the acquisition geometry shown in Fig. 1), and $\nabla^{2}$ is the Laplacian operator. The solution to this wave equation defines the forward modeling operator $f(\mathbf{x})$, which maps velocity models to synthetic seismic data.
FWI is formulated as a nonlinear optimization problem aimed at determining the velocity model $\mathbf{x}$ that minimizes the difference between the observed seismic data $\mathbf{u}_{\text{data}}$ and the synthetic data $f(\mathbf{x})$ generated by solving the wave equation numerically. This optimization problem can be expressed as
$\arg\min_{\mathbf{x}}\left[\frac{1}{2}\|\mathbf{u}_{\text{data}}-f_{\text{PDE}}(\mathbf{x})\|_{2}^{2}+\lambda R(\mathbf{x})\right],$
where $\|\cdot\|_{2}$ denotes the Euclidean norm, $R(\mathbf{x})$ represents a regularization term that incorporates prior information or enforces specific model characteristics, and $\lambda$ is a regularization parameter that balances data fidelity with regularization constraints.
The physics-driven nature of FWI is evident in the forward modeling operator $f(\mathbf{x})$, which requires solving the acoustic wave equation using numerical methods such as finite differences. In this process, synthetic seismic data is generated by simulating wave propagation from known source positions and recording the resulting wavefield at receiver locations, replicating the actual data acquisition geometry.
Despite its advantages, FWI presents significant computational and practical challenges due to its inherent nonlinearity and ill-posed nature. These challenges make FWI highly sensitive to noise and inaccuracies in the initial model. Noise in seismic data acquisition is inevitable and can severely degrade inversion performance. To mitigate these effects and improve the robustness of the inversion process, different regularization techniques are employed. These techniques integrate prior geological information and enforce model constraints, resulting in physically plausible velocity models.
Appendix S2 RED-DiffEq algorithm
Algorithm S1 RED-DiffEq for full waveform inversion.
0: Seismic data $\mathbf{\mathbf{u}_{\text{data}}}$, wave equation solver $f_{\text{PDE}}(\cdot)$, iterations $K$, noise schedule $\{\gamma(t)\}_{t=1}^{T}$, pre-trained diffusion model $\hat{\boldsymbol{\epsilon}}_{\theta}$, regularization parameter $\lambda$, step size $\eta$, $T=1000$
1: Initialize velocity model: $\mathbf{x}_{0}$
2:for $k=0,\ldots,K-1$ do
3:  $t\sim\mathcal{U}\{1:T\}$ {Sample timestep uniformly (integer)}
4:  $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ {Sample random Gaussian noise}
5:  $\mathbf{x}_{k,t}=\sqrt{\gamma(t)}\mathbf{x}_{k}+\sqrt{1-\gamma(t)}\boldsymbol{\epsilon}$ {Perturb velocity model at iteration kk}
6:  $\hat{\boldsymbol{\epsilon}}=\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)$ {Predict noise using pretrained diffusion model}
7:  $\mathcal{L}(\mathbf{x})=\|\mathbf{u}_{\text{data}}-f_{\text{PDE}}(\mathbf{x})\|_{2}^{2}+\lambda\mathbf{x}^{\top}(\text{sg}(\hat{\boldsymbol{\epsilon}})-\boldsymbol{\epsilon})$ {Calculate the loss, stop gradient (sg) on $\hat{\epsilon}$}
8:  $\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{k}-\eta\nabla_{\mathbf{x}}\mathcal{L}(\mathbf{x})|_{\mathbf{x}=\mathbf{x}_{k}}$ {Update via gradient descent (e.g., Adam)}
9:end for
10:return $\mathbf{x}_{K}$ {Final velocity model}
In our experiment, we adopt the Adam optimizer for implementing gradient descent. The noise schedule $\{\gamma(t)\}_{t=1}^{T}$ specifies the cumulative signal‐ratio at each diffusion step $t=1,\dots,T$. In diffusion models, the noise schedule dictates how noise is progressively added to the input data over $T$ time steps, transforming it into pure noise in the forward process. This schedule is critical for training the model to reverse the process and recover the original signal. In our implementation, we adopt a sigmoid‑based $\gamma$–schedule [68] (Algorithm S2 and Fig. S1).
Algorithm S2 Sigmoid $\gamma$ schedule.
0: Total steps $T=1000$, start S=3S=-3, end $E=3$, sharpness $\tau=1$
[…]
3:for $t=1$ to $T$ do
4:  $u_{t}\leftarrow\tfrac{t}{T}$
5:  $s_{t}\leftarrow\tfrac{u_{t}\,(E-S)+S}{\tau}$
6:  $\gamma_{t}\leftarrow\dfrac{\,v_{E}-\sigma(s_{t})\,}{\,v_{E}-v_{S}\,}$
7:end for
8:return $\{\gamma_{t}\}_{t=1}^{T}$
In diffusion models, the signal-to-noise ratio (SNR) at each step quantifies the relative strength of the remaining clean data versus the accumulated noise during the forward diffusion process. It is defined as $\mathrm{SNR}(t)=\gamma(t)/(1-\gamma(t))$, and thus $\gamma(t)$ is also called the cumulative signal-retention factor. As diffusion progresses, the model learns to predict increasingly noisy data, and the SNR directly influences the training dynamics. The sigmoid noise schedule causes a rapid transition from high-SNR (signal-dominant) to low-SNR (noise-dominant) regimes.
Figure S1: Sigmoid noise schedule. (Left) Cumulative signal‑retention factor $\gamma(t)$ and its complement $1-\gamma(t)$, illustrating the decay of clean signal versus accumulated noise. (Right) Signal‑to‑noise ratio $\gamma(t)/(1-\gamma(t))$, illustrating the transition from high‑SNR to low‑SNR regimes over time.
Appendix S3 Mathematical framework of RED-DiffEq
Throughout, let $\gamma(t)\in(0,1]$ denote the variance-preserving (VP) noise schedule at step $t\in\{1,\dots,T\}$. Let $p_{\text{data}}$ be the density of the clean prior distribution.
S3.1 Score function by DDPM
The marginal score is defined as
$s_{\text{data},t}(\mathbf{z})\coloneq\nabla_{\mathbf{z}}\log p_{\text{data},t}(\mathbf{z}).$
Under the DDPM framework [49], for a network $\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{z},t)$ trained with the standard $\epsilon$-prediction objective, we have the approximation
$s_{\text{data},t}(\mathbf{z})\;\approx\;-\,\frac{\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{z},t)}{\sqrt{1-\gamma(t)}}.$ (S1)
S3.2 Posterior mean for the VP process
Let $\mathbf{y}=\mathbf{x}+\sigma\,\boldsymbol{\epsilon}$ with $\boldsymbol{\epsilon}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$ and the noise variance $\sigma^{2}$. Tweedie’s formula [69] states
$\mathbb{E}\!\big[\mathbf{x}\,\big|\,\mathbf{y}\big]\;=\;\mathbf{y}+\sigma^{2}\nabla_{\mathbf{y}}\log p(\mathbf{y}).$
Under the VP channel $\mathbf{x}_{t}=\sqrt{\gamma(t)}\,\mathbf{x}+\sqrt{1-\gamma(t)}\,\boldsymbol{\epsilon}$, this gives
$\mathbb{E}\!\big[\mathbf{x}\,\big|\,\mathbf{x}_{t}\big]$ $=\;\frac{1}{\sqrt{\gamma(t)}}\Big[\mathbf{x}_{t}+\big(1-\gamma(t)\big)\,\nabla_{\mathbf{x}_{t}}\log p_{\text{data},t}(\mathbf{x}_{t})\Big]$ (S2)
$=\;\frac{1}{\sqrt{\gamma(t)}}\Big[\mathbf{x}_{t}+\big(1-\gamma(t)\big)\,s_{\text{data},t}(\mathbf{x}_{t})\Big].$
S3.3 RED estimator
Let $\mathbf{x}\in\mathbb{R}^{n}$ be the current solution during inversion. Using Eqs. (S1) and (S2), a practical denoiser is
$\mathcal{D}_{\theta}(\mathbf{x};t,\boldsymbol{\epsilon})\;=\;\frac{1}{\sqrt{\gamma(t)}}\Big(\mathbf{x}_{t}\;-\;\sqrt{1-\gamma(t)}\,\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\Big),\quad\mathbf{x}_{t}=\sqrt{\gamma(t)}\,\mathbf{x}+\sqrt{1-\gamma(t)}\,\boldsymbol{\epsilon}.$
Then for the regularization $R(\mathbf{x})=\mathbb{E}_{t,\boldsymbol{\epsilon}}\big[\mathbf{x}^{\top}\!\bigl(\mathbf{x}-\mathcal{D}_{\theta}(\mathbf{x};t,\boldsymbol{\epsilon})\bigr)\big]$, we estimate it with one Monte Carlo sample to get
$\mathbf{x}^{\!\top}\!\big(\mathbf{x}-\mathcal{D}_{\theta}(\mathbf{x};t,\boldsymbol{\epsilon})\big)$ $=\mathbf{x}^{\!\top}\!\left[\mathbf{x}-\frac{\sqrt{\gamma(t)}\,\mathbf{x}+\sqrt{1-\gamma(t)}\,\boldsymbol{\epsilon}-\sqrt{1-\gamma(t)}\,\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\gamma(t)}}\right]$
$=\sqrt{\frac{1-\gamma(t)}{\gamma(t)}}\;\mathbf{x}^{\!\top}\!\Big(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)-\boldsymbol{\epsilon}\Big)$
$\coloneq w(t)\,\mathbf{x}^{\!\top}\!\Big(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)-\boldsymbol{\epsilon}\Big).$
Appendix S4 Experiment details and baseline methods
The diffusion model was trained on the OpenFWI dataset [62], encompassing FF-B, FV-B, CV-B, and CF-B velocity model families, using a U-Net architecture with ResBlock and Attention Block components (Fig. 2a). The model architecture incorporated 1,000 diffusion steps with a sigmoidal noise schedule (start = 3-3, end = 3, $\tau=1$) in Alg. S2, 64 base channels with progressive channel multipliers (1, 2, 4, 8), and 4 attention heads. We trained the model for 400,000 iterations using a batch size of 32 and a learning rate of 0.0002.
To ensure fair comparison, hyperparameters were kept consistent across all experiments. The optimization process used a learning rate of 0.03 with a cosine annealing scheduler for 300 iterations. Regularization coefficients were tuned through empirical validation: 0.01 for both Tikhonov and Total Variation methods, and 0.7 for RED-DiffEq. These values were selected to achieve a balance between quantitative accuracy and qualitative fidelity, ensuring the regularization terms effectively constrained the solution space while avoiding common artifacts (such as the staircase effect) that compromise geological interpretation.
S4.1 Tikhonov regularization
Tikhonov regularization [17] is a widely used technique that promotes smoothness in the velocity model by penalizing large variations in parameter values. We apply the first-order Tikhonov regularization as
$R_{\text{Tikhonov}}(\mathbf{x})=\frac{1}{N}\sum_{i,j}\left((\mathbf{x}_{i+1,j}-\mathbf{x}_{i,j})^{2}+(\mathbf{x}_{i,j+1}-\mathbf{x}_{i,j})^{2}\right),$
where $\mathbf{x}_{i,j}$ represents the velocity value at the discrete position $(i,j)$, and $N$ is the total number of grid points. $R_{\text{Tikhonov}}(\mathbf{x})$ discourages abrupt changes in the velocity field, resulting in a smoother solution. This method is advantageous when the true subsurface structures are expected to vary smoothly.
S4.2 Total variation regularization
Total Variation (TV) regularization [15] is a powerful technique for preserving sharp interfaces and discontinuities in the velocity model, which are critical for representing geological boundaries such as faults. In our implementation, we apply the anisotropic TV regularization as
$R_{\text{TV}}(\mathbf{x})=\frac{1}{N}\sum_{i,j}\left(|\mathbf{x}_{i+1,j}-\mathbf{x}_{i,j}|+|\mathbf{x}_{i,j+1}-\mathbf{x}_{i,j}|\right),$
where $\mathbf{x}_{i,j}$ represents the velocity value at the discrete position $(i,j)$, and $N$ is the total number of grid points. However, a common drawback of TV regularization is the potential introduction of staircase artifacts, where smooth gradients are approximated by discrete, flat regions separated by abrupt transitions.
Appendix S5 Statistical test of performance for missing traces
To more rigorously assess the impact of missing traces on each method, we conducted a two-way ANOVA for three evaluation metrics: RMSE, MAE, and SSIM. The $p$-values associated with the number of missing traces (5–20) were 0.456, 0.568, and 0.624 for RMSE, MAE, and SSIM, respectively—well above the 0.05 significance threshold—indicating no statistically significant effect of missing traces alone.